100 grams of ice cream equals 15 grams of fat, how many grams of fat are in 260 grams of ice cream?
Answer by stanbon(75887)

100 grams of ice cream equals 15 grams of fat, how many grams of fat are in 260 grams of ice cream?
-------
Proportion:
fat/grams = fat/grams
----
x/260 = 15/100
---
x = 260(3/20)
---
x = 39 grams of fat
=========================
Cheers,
Stan H.
==========